given 2 sets A and B, use membership table to show that (A-B)∪(B-A)=(A∪B) -(A∩B)
1
Expert's answer
2020-01-10T08:24:52-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 A & B & A-B & B-A & (A-B\cup (B-A) \\ \hline
 1 & 1 & 0 & 0 & 0 \\ \hline
 1 & 0 & 1 & 0 & 1\\ \hline
 0 & 1 & 0 & 1 & 1\\ \hline
 0 & 0 & 0 & 0 & 0
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 A\cup B & A\cap B & (A\cup B)-(A\cap B) \\ \hline
 1 & 1 & 0 \\ \hline
 1 & 0 & 1 \\ \hline
 1 & 0 & 1 \\ \hline
 0 & 0 & 0 
\end{array}"

Hence


"(A-B)\cup(B-A)=(A\cup B)-(A\cap B)"